OpenStudy (dinorap1):

Please help! Will give medal to best answer! 6. Suppose a parabola has a vertex of (5,-3) and passes through the point (6,1). Write the equation of the parabola in vertex form. A. y=(x-5)^2-3 B. y=4(x-5)^2-3 C. y=4(x+5)^2-3 D. y=4(x-5)^2+3

OpenStudy (campbell_st):

well knowing the vertex you can write it as \[y = a(x - h)^2 + k\] where (h, k) is the vertex. so substitute your vertex. to find the value of a, substitute the point (6, 1) into the vertex equation to get thew value of a. then you'll have your equation

OpenStudy (dinorap1):

Oh, I see... so that's how you're supposed to do it. Okay, I'll try it and see what I get.

OpenStudy (campbell_st):

glad to help

OpenStudy (dinorap1):

But I've input it into my graphing calculator, and the only answers that might be right are A or B. Is that right or wrong?

OpenStudy (dinorap1):

If I knew how to find certain coordinates, I would already know the answer by now...

OpenStudy (campbell_st):

ok... so the vertex form after putting the vertex in is \[y =a(x - 5)^2 - 3\] now substitute x = 6 and y = 1 to find a

OpenStudy (dinorap1):

So when you plug that in, it becomes: 1=a(6-5)^2-3 Do you distribute a first, or do you distribute the ^2 first?

OpenStudy (campbell_st):

no just go \[1 = a \times (1)^2 -3\] whcih becomes 1 = a - 3 solve for a

OpenStudy (dinorap1):

Sorry, I get confused whenever there's an exponent hanging around... Okay, that makes sense. So: a=4

OpenStudy (campbell_st):

thats correct so you should eb able to identify the correct equation now

OpenStudy (dinorap1):

It's B, right? Thanks for the help by the way! :)

OpenStudy (campbell_st):

that's correct